JamesTDG:

The instruction did not do a good job teaching me, while I have somehow guessed the answer, I would prefer it if somebody could teach me how to solve this. simplify sqrt-50 A | 5 sqrt2 B | 5i sqrt2 C | -5 sqrt2 D | -5i sqrt2

Serenity1jacksonsl:

Ok um do you know how to divide your cordinates

Vocaloid:

remember your rules for expanding radicals sqrt(a*b) = sqrt(a) * sqrt(b) using this logic, we can re-write sqrt(-50) as sqrt(-1) * sqrt(50) sqrt(-1) is simply i by definition, so we have i * sqrt(50) after that, you can re-write sqrt(50) as sqrt(25*2) using the same logic as before, and simplify

surjithayer:

\[\sqrt{a}\times \sqrt{b}=\sqrt{ab}\] if atleast one of a or b >0